OpenStudy (anonymous):

Find d2y/dx2 in terms of x and y if (x^2)-(y^2)=16

OpenStudy (turingtest):

differentiate implicitly and solve for dy/dx show what you get

OpenStudy (anonymous):

i got for y'=2x-2y and for y''=0

OpenStudy (anonymous):

that just doesnt look right to me

OpenStudy (turingtest):

d/dx(x^2)+d/dx(y^2)=d/dx(5) 2x+2y(dy/dx)=0 dy/dx=-x/y

OpenStudy (turingtest):

now differentiate implicitly again whenever you get dy/dx in what you are doing you can sub in -x/y

OpenStudy (anonymous):

how did you get 5

OpenStudy (turingtest):

or 16 whatever it's a constant so either way it is zero upon taking the derivative

OpenStudy (turingtest):

d/dx(x^2)+d/dx(y^2)=d/dx(16) 2x+2y(dy/dx)=0 dy/dx=-x/y

OpenStudy (turingtest):

doesn't change anything

OpenStudy (anonymous):

ok i see it now

OpenStudy (anonymous):

so how do i get the second prime from that? i dont see

OpenStudy (turingtest):

d/dx(dy/dx)=d/dx(-x/y) you can use the quotient rule

OpenStudy (anonymous):

ok i got -16/y^3

OpenStudy (turingtest):

where did -16 come from?

OpenStudy (anonymous):

idk thats what i got when i put it in wolfram look http://www.wolframalpha.com/input/?i=%28x%5E2%29-%28y%5E2%29%3D16

OpenStudy (turingtest):

I see nothing about the second derivative in that link

OpenStudy (anonymous):

you have to press more next to implicit derivatives

OpenStudy (turingtest):

oh ic

OpenStudy (turingtest):

well it might help to derive that, plus I would imagine they want that in terms of x anyway

OpenStudy (anonymous):

sorry im just really bad with derivatives

OpenStudy (anonymous):

?

OpenStudy (turingtest):

\[x^2+y^2=16\]\[2x+2yy'=0\implies y'=-\frac xy\]\[y''=-\frac d{dx}\left(\frac xy\right)=-{y\frac d{dx}(x)-x\frac d{dx}(y)\over y^2}\]\[y''=-{y-x(-\frac xy)\over y^2}=-\frac{y^2+x^2}{y^3}=-\frac{16}{y^3}\]

OpenStudy (anonymous):

now i see it !!! is that chain rule?

OpenStudy (turingtest):

yes, exactly chain rule on y(x) d/dx(f(y))=f'(y)(dy/dx) <- chain rule

OpenStudy (anonymous):

i didn't know that you could use chain rule with fractions that way

OpenStudy (turingtest):

the quotient rule is really just the product rule + chain rule:\[(\frac uv)'=(uv^{-1})'=u'v^{-1}-uv^{-2}v'=\frac{u'v-v'u}{v^2}\]

OpenStudy (anonymous):

i see!! thank you so much